Payments of $ 670 are being made at the end of each month for 5 years at an interest of 8% compounded monthly. Calculate the Present Value.
Answer #1
Present value = Monthly payment * Present value of annuity of 1
= $       670.00 * 49.31843
= $ 33,043.35
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.006667)^-60)/0.006667 i = 8%/12 = 0.006667
= 49.3184333 n = 5*12 = 60